OpenStudy (anonymous):

The area of a square garden is 128m squae. How long is the diagonal?

OpenStudy (anonymous):

Find the square root first. Then use the pythagorean theorum to find the diagonal.

OpenStudy (anonymous):

128=S^2 \[s=\sqrt{128}\]

OpenStudy (anonymous):

\[D^2=\sqrt{128}^2+\sqrt{128}^2\]

OpenStudy (anonymous):

\[D=\sqrt{256}\]

OpenStudy (anonymous):

D=16

OpenStudy (id21):

\[Area=a^2\] \[a^2=128\] \[a=8\sqrt{2}\] diagonal of square \[= a \sqrt 2 = 8 \sqrt{2\times2}=16\]